OpenStudy (juanilet2000):

If equation of Circle B is x^2 + y^2 - 4x +12y - 7= 0, find the center and the radius of circle B. Center = (_,_), Radius =

OpenStudy (juanilet2000):

@3mar

OpenStudy (3mar):

Well, I am here. That is with my pleasure to help you!

OpenStudy (juanilet2000):

Thank you!

OpenStudy (3mar):

Not at all! Are you familiar with circle equation?

OpenStudy (juanilet2000):

No I didn't know the formula, thanks for telling me

OpenStudy (3mar):

So can you deal with it?

OpenStudy (juanilet2000):

Eh, its confusing because I'm not sure whether I'm suppose to plug in the numbers

OpenStudy (3mar):

Are you familiar with completing a square?

OpenStudy (juanilet2000):

Yes.

OpenStudy (3mar):

Great! Can you do that for this x and y in that relation (x^2 + y^2 - 4x +12y - 7= 0)??

OpenStudy (3mar):

x^2 - 4x ?????? y^2 + 2y ??????

OpenStudy (juanilet2000):

Yeah, sorry for the long reply. I did (x^2 - 4x) (y^2 + 2y) - 7 = 0

OpenStudy (3mar):

No No No No Not multiplication! I said : complete the square!!!

OpenStudy (juanilet2000):

I didn't multiply. Now I'm lost

OpenStudy (mathmale):

Looks as tho' Juanilet might benefit from your going thru an example of completing the square.

OpenStudy (3mar):

Look! \[\Large x^2 + y^2 - 4x +12y - 7= 0\] \[\Large (x^2-4x) + (y^2+12y) - 7= 0\] \[\Large (x^2-4x\color{red}{+4-4}) + (y^2+12y\color{red}{+36-36}) - 7= 0\] \[\Large ([x^2-4x+4]-4) + ([y^2+12y+36]-36) - 7= 0\] \[\Large ([x-2]^2-4) + ([y+6]^2-36) - 7= 0\] \[\Large [x-2]^2 + [y+6]^2-36 - 7-4= 0\] \[\Large [x-2]^2 + [y+6]^2-47= 0\] \[\Large [x-2]^2 + [y+6]^2= 47\] I hope that helps!

OpenStudy (juanilet2000):

Okay thank you!

OpenStudy (3mar):

Are you satisfied?

OpenStudy (juanilet2000):

A little I'm just dumb.

OpenStudy (3mar):

Why? I need you 100 % sure!

OpenStudy (3mar):

about completing the square or center and radius of the circle?

OpenStudy (3mar):

@Juanilet2000 Do you follow?

OpenStudy (juanilet2000):

I understand the completing the square I'm just confused on the radius of the circle.

OpenStudy (3mar):

 \[\Huge\color{green}ツ\]